Circular No. 5622

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET BREWINGTON (1992p) Further precise positions have been reported as follows:

1992 UT α_{2000} δ_{2000} m_{1} Observer
Sept. 1.50469 7^{h}24^{m}42.^{s}63 +36^{o}44′37′′.3  Tatum
1.79861 7 25 30.41 +36 43 54.8 14  Otomo
2.47274 7 27 19.22 +36 42 08.1  Tatum
4.79601 7 33 30.24 +36 35 51.6 14  Seki
4.81181 7 33 32.70 +36 35 49.5  _{”}
25.74271 8 23 31.70 +35 17 16.4 16  _{”}
25.75035 8 23 32.56 +35 17 14.4  _{”}
26.77535 8 25 42.95 +35 12 50.6 16.5  _{”}
26.78056 8 25 43.71 +35 12 50.1  _{”}

J. B. Tatum (University of Victoria). 0.25-m Schmidt. S. Otomo (Otomo, Japan). 0.25-m f/3.4 Schmidt. T. Seki (Geisei). 0.60-m f/3.5 reflector.

On communicating the late-September observations, S. Nakano, Sumoto, Japan, remarks that the comet is evidently of short period. The following improved orbital elements have been determined by the undersigned from 18 observations Aug. 28-Sept. 26:

T = 1992 June 4.050 TT  ω = 45.^{o}223
e = 0.62962  Ω = 342.919 }2000.0
q = 1.56038 AU  i = 18.096

a = 4.21290 AU n^{o} = 0.113981 P = 8.65 years

1992TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 25 8^{h}21.^{m}93 +35^{o}20.^{′}5 2.185 1.963 63.^{o}9 27.^{o}3 16.6
Oct. 5 8 42.17 +34 37.5 2.150 2.024 69.2 27.5 16.7
15 8 59.83 +33 56.3 2.109 2.087 75.0 27.5 16.8
25 9 14.81 +33 20.6 2.062 2.151 81.4 27.2 16.9
Nov. 4 9 26.95 +32 53.6 2.010 2.217 88.4 26.6 17.0
14 9 36.07 +32 37.5 1.955 2.283 96.1 25.5 17.0
1992 QB_{1}

Further observations by D. Jewitt at Mauna Kea on Sept. 25 confirm the validity of the ephemeris on IAUC 5611. The orbit is still very indeterminate, but retrograde solutions can probably be precluded.

1992 September 27 Brian G. Marsden